OpenStudy (jaylelile):

Please help me! Will fan and medal! A candy merchant wants to mix peanuts selling at $2.20 per pound with cashews selling at $5.40 per pound to form 120 lbs of a mixed-nut blend that will sell for $3 per pound. What amount of each type of nut should be used?

OpenStudy (xariazs):

So if we start by naming our variables we get that x= peanuts and y=cashews x+y=120 2.20x+5.40y=120(3) --> 2.20x+5.40y=360 From here on you can do some substitution or elimination.

OpenStudy (xariazs):

The equations correspond respectively to the pounds and the price, so if you notice that first equation deals with the pounds, while the second equation deals with the money